# centertableaux,boxsize=.5em

### v1

@@
[…]
centertableaux,boxsize=.5em
1 Department of Mathematics, King’s College London, Strand, London WC2R 2LS, UK.
@@
Anomalies of Coset Non-Invertible Symmetries
Po-Shen Hsin1, Ryohei Kobayashi2, Carolyn Zhang3
@@ 1 Introduction
In recent years, generalization of group-like symmetry, called non-invertible symmetry, has developed rapidly with many examples in field theories and lattice models (see e.g. [1, 2, 3, 4, 5, 6, 7, 8, 9]). Non-invertible symmetry differs from group-like symmetry in that their fusion rules are not invertible. The symmetry operators do not obey a group multiplication law, and instead form a (higher) fusion category. While there is much progress in understanding mathematical properties of non-invertible symmetries, they often requires sophisticated techniques in fusion categories which could be challenging to compute explicitly.
In this work, we will study a class of non-invertible symmetries that come from cosets $G/K$ where $K\subset G$ is a not a normal subgroup of $G$, continuing our previous work [10] (see also [11, 12, 13, 14, 15, 16, 17, 18, 19] for other examples). The coset symmetry is obtained by starting with a theory with $G$ symmetry and gauging the possibly non-normal subgroup $K$. The coset symmetry defects can be constructed using a sandwich construction. Inside the sandwich we have a $G$ symmetry defect (possibly with ’t Hooft anomalies), and outside the sandwich the non-anomalous $K$ subgroup is gauged, possibly with a topological action, and the interfaces on the two sides correspond to Dirichlet boundary condition of $K$ gauge field (see Fig. 1).
In this definition, the coset symmetry comes with $\text{Rep}(K)$ symmetry– they are generated by the Wilson lines of $K$ gauge field. In particular, this means that although as cosets $G^{\prime}/K^{\prime}=G/K$ for $K^{\prime}=K/N,G^{\prime}=G/N$ for common normal subgroup $N$ of $G,K$, the symmetries $G/K$ and $G^{\prime}/K^{\prime}$ differ by $\text{Rep}(K),\text{Rep}(K^{\prime})$. The groups $G,K$ depend on the spectrum of topological lines – if they form $\text{Rep}(K)$, then the pair of groups are $(G,K\subset G)$ such that $G/K$ equals the coset. In particular, if the coset is a group and there are no topological line operators, then $K=1$. If the coset is not a group, then there must be topological line operators. We will discuss in more the detail the definition of coset symmetry in Sec. 2. In particular, we will introduce suitable twists for the symmetry that can change the fusion rules and higher group structures of the coset symmetry.
@@ 1.2 Summary of results
We show a general coset symmetry in $D$ spacetime dimensions can be described by the quadruple $(G,K,\omega_{D+1},\alpha_{D})$, where $K$ is a subgroup of $G$, $\omega_{D+1}$ is a $(D+1)$-cocycle for group $G$ with $U(1)$ coefficient, such that $\omega_{D+1}|_{K}=d\alpha_{D}$. This is a generalization of the group theoretical fusion category for $D=2$ [31] and the group theoretical fusion 2-category for $D=3$ [32] to generic spacetime dimensions. The twists have the following consequences on the algebraic structure of the coset symmetry:
@@
  • The twist $\alpha_{D}$ can affect the action of the coset symmetry: part of the symmetry may act trivially for different choices of $\alpha_{D}$.
@@
In addition, the presentation of the coset symmetry $(G,K,\omega_{D+1},\alpha_{D})$ is in general redundant: there can be two presentations describe the same symmetry. This can arise when the bulk TQFT admits two (or more) descriptions: one as $G$ gauge theory with topological action $\omega_{D+1}$, and another as $G^{\prime}\neq G$ gauge theory with topological action $\omega_{D+1}^{\prime}$.
@@
  • If a finite group $G=H\Join K$ is a bicrossed product of $K$ with another group $H$, then the coset symmetry in $D\geq 3$ can be realized in symmetric gapped phases. 111 For two subgroups $H,K$ of $G$, $G$ is the bicrossed product $G=H\Join K$ if and only if $G$ can be expressed as $G=HK$ and $H\cap K=\{\text{id}\}$. Equivalently, each group element $g\in G$ has a unique expression as $g=hk$ with $h\in H,k\in K$. Furthermore, if one can find a choice of the subgroup $H\subset G$ in the expression $G=H\Join K$ such that the twist $\omega_{D+1}$ restricted to $H$ becomes trivial, i.e., $[\omega_{D+1}|_{H}]=0$ in $H^{D+1}(BH,U(1))$, then the coset symmetry is anomaly free and can be realized in SPT phases.
@@
Any system with coset symmetry with anomalies or fractional responses must flow to nontrivial phases. We present examples of systems with fractional responses of coset symmetry that originate from nontrivial twists $\omega_{D+1},\alpha_{D}$. For example, the twisted coset symmetry in two 2+1D massless Dirac fermions coupled to $\mathbb{Z}_{2}$ gauge field can be enriched with twisted coset symmetry $(G=S_{3},K=\mathbb{Z}_{2},\omega_{4}=d\eta_{3}^{\prime}/2,\alpha_{3}=\eta_{3}/2)$ where $\eta_{3}^{\prime}$ and $\eta_{3}$ are the $\mathbb{Z}_{3}$ and $\mathbb{Z}_{2}$ generators of $H^{3}(BS_{3},U(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{3}$, respectively. Such twisted coset symmetry exhibits fractional response and cannot be realized in SPT phases.
@@ 2.1 Definition of twisted coset symmetry: $(G,K,\omega_{D+1},\alpha_{D})$
In this section, we will generalize the definition of coset symmetry in [10] for spacetime dimension $D$, into the following data $(G,K,\omega_{D+1},\alpha_{D})$:
@@
  • Topological action for $G$ in $D+1$ spacetime dimension, $\omega_{D+1}\in H^{D+1}(BG,U(1))$.
@@
    where $\omega_{D+1}|_{K}$ is the pullback under the inclusion $K\rightarrow G$222 A similar constraint is also imposed in the study of Cheshire strings in twisted gauge theories [34]. Different $\alpha_{D}$ are related by $\alpha_{D}\rightarrow\alpha_{D}+\eta_{D}$ with $\eta_{D}\in H^{D}(BK,U(1))$. When $\eta_{D}$ is obtained by restricting the $G$ SPT response $\eta^{\prime}\in H^{D}(BG,U(1))$ to $K\subset G$, $\eta_{D}$ represents invertible gauged SPT symmetry of the bulk $G$ gauge theory [35]. In this case, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ does not modify the symmetry category, since the gauged SPT defect $\eta_{D}$ induces the automorphism of the topological operators at the boundary. Meanwhile, when $\eta_{D}$ is not obtained by the restriction of $G$ SPT response to $K$, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ generally modifies the symmetry category of operators at the boundary, including its fusion rule. Nontrivial $\alpha_{D}$ that cannot be removed by $\eta_{D}$ represents a fractional response.
@@
Description using bulk TQFT
@@
The bulk $G$ gauge theory has higher group symmetry that depends on the bulk cocycle $\omega_{D+1}$ as studied in [35]. In the rest of the section, we will use the bulk description to study properties of coset symmetry via bulk-boundary correspondence. We will shortly see that the higher-group structure in the bulk $G$ gauge theory affects on the algebraic structure of the coset symmetry $(G,K,\omega_{D+1},\alpha_{D})$. Later in Sec. 3, we will use the bulk TQFT description to study anomalies of coset symmetries as obstructions to realization in symmetric, invertible phases.
@@ 2.2 Modified fusion rules from twist
The coset 0-form symmetry defects correspond to the magnetic defects in the bulk $G$ gauge theory. The pure magnetic defects are labeled by the conjugacy classes of $G$. When the gauge group is non-Abelian, the product of two conjugacy classes $[g_{1}],[g_{2}]$ give a direct sum of conjugacy classes $\oplus_{g}[g_{1}gg_{2}g^{-1}]$, and thus the fusion of magnetic fluxes is in general non-Abelian even in the absence of a topological twist $\omega_{D+1}$ for the $G$ gauge field.333 Note that in general, the fusion of pure magnetic fluxes can also include electric defects: these are the electric defects that become trivial when restricted to the common stabilizer subgroup of $G$ preserved by the fluxes. When the conjugacy class is in the center of $G$, the fusion of these fluxes is Abelian for $\omega_{D+1}=0$.
@@
As discussed in [35], the topological action $\omega_{D+1}$ of $G$ gauge theory can modify the fusion rule of the magnetic defects. This is because the magnetic defect of holonomy $g\in G$ extended in $(D-1)$ dimensions is attached to the additional gauged SPT defect in $D$ dimensions given by the slant product $i_{g}\omega_{D+1}$, forming the bulk-boundary operator (see Figure 2). In other words, the magnetic defect carries the non-invertible theory with anomalous $G$ symmetry which is the boundary of SPT $i_{g}\omega_{D+1}$, and this makes additional contributions to their non-invertible fusion rules. For the magnetic defect of holonomy $g\in G$, the gauged SPT defects attached to it is expressed as
$i_{g}\omega_{D+1}=i_{g}^{A}\omega_{D+1}+\frac{1}{|\omega_{D+1}|}d\left(i_{g}^{B}\omega_{D+1}\right)~{}.$ (2.5)
where $|\omega_{D+1}|$ is the order of $[\omega_{D+1}]$ in $H^{D+1}(BG,U(1))$.
@@
The contribution from $i^{A}_{g}$ modifies the fusion rules of magnetic defects, while $i_{g}^{B}$ modifies the junction of magnetic defects [35] (see Figure 3). As a consequence, the fusion rules of the coset symmetry on the gapped boundary of the bulk $G$ gauge theory is also modified. The boundary condition breaks $G$ to subgroup $K$, and the boundary action for $K$ is $\alpha_{D}$. Below, we describe the fusion rule of coset symmetries realized by the topological defects of the boundary.
@@
$V_{g}\times V_{g^{\prime}}=V_{g+g^{\prime}}\frac{1}{{\cal N}}\left(\sum_{\lambda\in Z(G)}{\cal W}_{i_{\lambda}i_{g}\omega^{(D)}}\right)\left(\sum_{\lambda^{\prime}\in Z(G)}{\cal W}_{i_{\lambda^{\prime}}i_{g^{\prime}}\omega^{(D)}}\right)\big{/}\{{\cal W}_{i_{\lambda^{\prime\prime}}i_{g+g^{\prime}}(\omega^{(D)})}:\lambda^{\prime\prime}\in Z(G)\}~{},$ (2.6)
@@
$V_{g}\times{\cal W}_{i_{\lambda}i_{g}\omega^{(D)}}=V_{g}~{},\quad\forall\lambda\in Z(G)~{},$ (2.7)
@@
Reflecting such fusion rules of the magnetic defects in the bulk, the fusion rules of coset symmetry on the boundary are modified if $\omega_{D+1}$ satisfies $\omega_{D+1}|_{K}=d\alpha$ and also $[i_{g}^{A}\omega_{D+1}]_{K}\neq 0$. In other words, although the restriction of $\omega_{D+1}$ to subgroup $K$ is exact, it is no longer exact if we first take the slant product of $\omega_{D+1}$ with respect to $G$.
@@
The property that the coset symmetry is attached to a gauged SPT defect implies that the coset symmetry becomes “more non-invertible” in the presence of twists: for magnetic defects with center holonomy, there are non-invertible fusion rule obtained by replacing $i_{\lambda}i_{g}\omega^{(D)}$ in (2.6), (2.7) to $[i_{\lambda}i_{g}\omega^{(D)}]_{K}$.
@@
Let us briefly recall how the slant product $i^{B}$ modifies the fusion rule of magnetic defects in the bulk $G$ gauge theory. The slant product $i^{B}$ modifies the $(D-2)$-dimensional junction of the magnetic defects by attaching the end of the gauged SPT defects in $(D-1)$ dimensions. In other words, the fusion algebra of the magnetic defects is extended by the gauged SPT defects.
@@
$Z_{\omega}(G)\equiv\{g\in Z(G):\quad[i^{A}_{g}\omega^{(D)}]=0\in H^{D-1}(BG,U(1))\}~{}.$ (2.8)
@@
$1\rightarrow H^{D-2}(BG,U(1))\rightarrow{\cal G}^{(1)}\rightarrow Z_{\omega}(G)\rightarrow 1~{},$ (2.9)
which is characterized by the second cohomology $\Omega$ [35]
$\Omega(g,g^{\prime})=\frac{1}{|\omega^{(D)}|}\left(i_{g}^{B}\omega^{(D)}+i_{g^{\prime}}^{B}\omega^{(D)}-i_{g+g^{\prime}}^{B}\omega^{(D)}\right)~{}.$ (2.10)
If $\Omega$ is nontrivial when restricted to $K$, i.e., $\Omega(g,g^{\prime})|_{K}$ is nontrivial in cohomology $H^{2}(Z_{\omega}(G),H^{D-2}(BK,U(1)))$, then the coset 0-form symmetry at the boundary also has the central extension,
$1\rightarrow H^{D-2}(BK,U(1))\rightarrow{\cal G}^{(0)}_{G/K}\rightarrow Z_{\omega}(G)\rightarrow 1~{}.$ (2.11)

### v3

@@
1 Department of Mathematics, King’s College London, Strand, London WC2R 2LS, UK
@@
Anomalies of global symmetries provide important information on the quantum dynamics. We show the dynamical constraints can be organized into three classes: genuine anomalies, fractional topological responses, and integer responses that can be realized in symmetry-protected topological (SPT) phases. Coset symmetry can be present in many physical systems including quantum spin liquids, and the coset symmetry can be a non-invertible symmetry. We introduce twists in coset symmetries, which modify the fusion rules and the generalized Frobenius-Schur indicators. We call such coset symmetries twisted coset symmetries, and they are labeled by the quadruple $(G,K,\omega_{D+1},\alpha_{D})$ in $D$ spacetime dimensions where $G$ is a group and $K\subset G$ is a discrete subgroup, $\omega_{D+1}$ is a $(D+1)$-cocycle for group $G$, and $\alpha_{D}$ is a $D$-cochain for group $K$. We present several examples with twisted coset symmetries using lattice models and field theory, including both gapped and gapless systems (such as gapless symmetry-protected topological phases). We investigate the anomalies of general twisted coset symmetry, which presents obstructions to realizing the coset symmetry in (gapped) symmetry-protected topological phases. We show that finite coset symmetry $G/K$ becomes anomalous when $G$ cannot be expressed as the bicrossed product $G=H\Join K$, and such anomalous coset symmetry leads to symmetry-enforced gaplessness in generic spacetime dimensions. We illustrate examples of anomalous coset symmetries with $A_{5}/\mathbb{Z}_{2}$ symmetry, with realizations in lattice models.
November 5, 2025
@@ 1 Introduction
In recent years, generalization of group-like symmetry, called non-invertible symmetry, has developed rapidly with many examples in field theories and lattice models (see e.g. [1, 2, 3, 4, 5, 6, 7, 8, 9]). Non-invertible symmetry differs from group-like symmetry in that their fusion rules are not invertible. The symmetry operators do not obey a group multiplication law, and instead form a (higher) fusion category. While there is much progress in understanding mathematical properties of non-invertible symmetries, they often require sophisticated techniques in fusion categories which could be challenging to compute explicitly.
In this work, we will study a class of non-invertible symmetries that come from cosets $G/K$ where $K\subset G$ is a not necessarily a normal subgroup of $G$, continuing our previous work [10] (see also [11, 12, 13, 14, 15, 16, 17, 18, 19] for other examples). The coset symmetry is obtained by starting with a theory with $G$ symmetry and gauging the possibly non-normal subgroup $K$. The coset symmetry defects can be constructed using a sandwich construction. Inside the sandwich we have a $G$ symmetry defect (possibly with ’t Hooft anomalies), and outside the sandwich the non-anomalous $K$ subgroup is gauged, possibly with a topological action, and the interfaces on the two sides correspond to Dirichlet boundary condition of $K$ gauge field (see Fig. 1).
In this definition, the coset symmetry comes with $\text{Rep}(K)$ symmetry– they are generated by the Wilson lines of $K$ gauge field. In particular, this means that although as cosets $G^{\prime}/K^{\prime}=G/K$ for $K^{\prime}=K/N,G^{\prime}=G/N$ for common normal subgroup $N$ of $G,K$, the symmetries $G/K$ and $G^{\prime}/K^{\prime}$ differ by $\text{Rep}(K),\text{Rep}(K^{\prime})$. The groups $G,K$ depend on the spectrum of topological lines – if they form $\text{Rep}(K)$, then the pair of groups are $(G,K\subset G)$ such that $G/K$ equals the coset. In particular, if the coset is a group and there are no topological line operators, then $K=1$. If the coset is not a group, then there must be topological line operators. We will discuss more in the detail the definition of coset symmetry in Sec. 2. In particular, we will introduce suitable twists for the symmetry that can change the fusion rules and higher group structures of the coset symmetry.
@@ 1.2 Summary of results
We show a general coset symmetry in $D$ spacetime dimensions can be described by the quadruple $(G,K,\omega_{D+1},\alpha_{D})$, where $K$ is a subgroup of $G$, $\omega_{D+1}$ is a $(D+1)$-cocycle for group $G$ with $U(1)$ coefficient, such that $\omega_{D+1}|_{K}=d\alpha_{D}$. Such symmetry is obtained by starting with invertible $G$ symmetry with an ’t Hooft anomaly $\omega_{D+1}$, and then gauging a $K$ subgroup symmetry with the twist $\omega_{D}$. This is a generalization of the group theoretical fusion category for $D=2$ [31] and the group theoretical fusion 2-category for $D=3$ [32] to generic spacetime dimensions. The twists have the following consequences on the algebraic structure of the coset symmetry:
@@
  • The twist $\alpha_{D}$ can modify the fusion rules of the symmetry.
@@
In our discussion, we include the fractional topological response as part of the symmetry data similar to how systems with Lieb-Schultz-Mattis (LSM) anomalies from translation and filling constraints, one also has to specify the filling [33, 34]. The filling is not a piece of data from the symmetry algebra and indeed does not appear in the symTFT construction, but can still impose important constraints on dynamics. It would be interesting to understand how to incorporate this data into the symTFT description.
@@
In addition, the presentation of the coset symmetry $(G,K,\omega_{D+1},\alpha_{D})$ is in general redundant: there can be two presentations describe the same symmetry. For instance, this can arise when the bulk TQFT admits two (or more) descriptions: one as $G$ gauge theory with topological action $\omega_{D+1}$, and another as $G^{\prime}\neq G$ gauge theory with topological action $\omega_{D+1}^{\prime}$.
@@
  • If a finite group $G=H\Join K$ is a bicrossed product of $K$ with another group $H$, then the coset symmetry in $D\geq 3$ can be realized in symmetric gapped phases.  [^1] For two subgroups $H,K$ of $G$, $G$ is the bicrossed product $G=H\Join K$ if and only if $G$ can be expressed as $G=HK$ and $H\cap K=\{\text{id}\}$. Equivalently, each group element $g\in G$ has a unique expression as $g=hk$ with $h\in H,k\in K$. Furthermore, if one can find a choice of the subgroup $H\subset G$ in the expression $G=H\Join K$ such that the twist $\omega_{D+1}$ restricted to $H$ becomes trivial, i.e., $\omega_{D+1}|_{H}]=0$ in $H^{D+1}(BH,U(1))$, then the coset symmetry is anomaly free and can be realized in SPT phases.
@@
Any system with coset symmetry with anomalies or fractional responses must flow to nontrivial phases. We present examples of systems with fractional responses of coset symmetry that originate from nontrivial twists $\omega_{D+1},\alpha_{D}$. For example, the twisted coset symmetry in two 2+1D massless Dirac fermions coupled to $\mathbb{Z}_{2}$ gauge field can be enriched with twisted coset symmetry $O(2)/\mathbb{Z}_{2}$, with $\omega$ being the theta term corresponds to the $O(2)_{1/2,1/2}$ fractional Chern-Simons term and $\alpha=\frac{1}{2}\eta$ being half of the minimal $\mathbb{Z}_{2}$ Chern-Simons term, where the notation is the same as [35]. Such twisted coset symmetry exhibits fractional response and cannot be realized in SPT phases.
@@ 2.1 Definition of twisted coset symmetry: $(G,K,\omega_{D+1},\alpha_{D})$
In this section, we will generalize the definition of coset symmetry in [10] for spacetime dimension $D$, into the following data $(G,K,\omega_{D+1},\alpha_{D})$ that describes gauging a subgroup $K\subset G$ in invertible $G$ symmetry:
@@
  • Topological action for $G$ in $D+1$ spacetime dimension, $\omega_{D+1}\in H^{D+1}(BG,U(1))$. This represents that $G$ symmetry has an ’t Hooft anomaly characterized by bulk SPT response $\omega_{D+1}$.
@@
    where $\omega_{D+1}|_{K}$ is the pullback under the inclusion $K\rightarrow G$. [^2] A similar constraint is also imposed in the study of Cheshire strings in twisted gauge theories [37]. Different $\alpha_{D}$ are related by $\alpha_{D}\rightarrow\alpha_{D}+\eta_{D}$ with $\eta_{D}\in Z^{D}(BK,U(1))$. This represents that an anomaly free subgroup $K\subset G$ is gauged with the twist valued in $\alpha_{D}$, and the twist can be shifted by $D$-dimensional SPT response $\eta_{D}$. Such a shift in $\alpha_{D}$ may or may not change the symmetry category depending on the cohomology class $\eta_{D}]$, as we will see in Sec. 2.1.1.
@@
2.1.1 Description using bulk TQFT
@@
As we have seen above, $\alpha_{D}$ can be shifted by $\alpha_{D}\to\alpha_{D}+\eta_{D}$ with an element $\eta_{D}\in Z^{D}(BG,U(1))$. When $\eta_{D}]\in H^{D}(BG,U(1))$ is obtained by restricting the $G$ SPT response $\eta^{\prime}]\in H^{D}(BG,U(1))$ to $K\subset G$, the shift by $\eta_{D}$ represents the action of invertible global symmetry (called gauged SPT symmetry) of the bulk $G$ gauge theory on the boundary [38]. In this case, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ does not modify the symmetry category, since the gauged SPT defect $\eta_{D}$ induces the automorphism of the topological operators at the boundary. Meanwhile, when $\eta_{D}$ is not obtained by the restriction of $G$ SPT response to $K$, the shift $\alpha_{D}\to\alpha_{D}+\eta_{D}$ generally modifies the symmetry category of operators at the boundary, including its fusion rule. Nontrivial $\alpha_{D}$ that cannot be removed by $\eta_{D}$ represents a fractional response. The effect of $\alpha_{D}$ on the symmetry category structure will be discussed in detail in Sec. 2.3.
The bulk $G$ gauge theory has higher group symmetry that depends on the bulk cocycle $\omega_{D+1}$ as studied in [38]. In the rest of the section, we will use the bulk description to study properties of coset symmetry via bulk-boundary correspondence. We will shortly see that the higher-group structure in the bulk $G$ gauge theory affects the algebraic structure of the coset symmetry $(G,K,\omega_{D+1},\alpha_{D})$. Later in Sec. 3, we will use the bulk TQFT description to study anomalies of coset symmetries as obstructions to realization in symmetric, invertible phases.
@@ 2.2 Modified fusion rules from twist
The coset 0-form symmetry defects correspond to the magnetic defects in the bulk $G$ gauge theory. The pure magnetic defects are labeled by the conjugacy classes of $G$. When the gauge group is non-Abelian, the product of two conjugacy classes $g_{1}],[g_{2}]$ give a direct sum of conjugacy classes $\oplus_{g}[g_{1}gg_{2}g^{-1}]$, and thus the fusion of magnetic fluxes is in general non-Abelian even in the absence of a topological twist $\omega_{D+1}$ for the $G$ gauge field. [^3] Note that in general, the fusion of pure magnetic fluxes can also include electric defects: these are the electric defects that become trivial when restricted to the common stabilizer subgroup of $G$ preserved by the fluxes. When the conjugacy class is in the center of $G$, the fusion of these fluxes is Abelian for $\omega_{D+1}=0$.
@@
As discussed in [38], the topological action $\omega_{D+1}$ of $G$ gauge theory can modify the fusion rule of the magnetic defects. This is because the magnetic defect of holonomy $g\in G$ extended in (D1)(D-1) dimensions is attached to the additional gauged SPT defect in $D$ dimensions given by the slant product $i_{g}\omega_{D+1}$. The attached gauged SPT can be decomposed into two parts: one part, called $i_{g}^{A}\omega_{D+1}$ can make the magnetic defects obey non-invertible fusion rules. The other part, called, $i_{g}^{B}\omega_{D+1}$, modify the junction of magentic defects. They are explicitly expressed as follows.
$i_{g}\omega_{D+1}=i_{g}^{A}\omega_{D+1}+\frac{1}{|\omega_{D+1}|}d\left(i_{g}^{B}\omega_{D+1}\right)~,$ (2.5)
where $|\omega_{D+1}|$ is the order of $\omega_{D+1}]$ in $H^{D+1}(BG,U(1))$.
The above operations $i_{g}^{A},i_{g}^{B}$ are defined as follows. Suppose that $\omega_{D+1}]$ has the order of $k=|\omega_{D+1}|$ in $H^{D+1}(BG,U(1))$, then one can fix a cocycle representative which takes value in $2\pi\mathbb{Z}/k$ for any $(D+2)$ group elements:
$\omega_{D+1}={2\pi\over k}\left(\omega_{D+1}\right)_{k}\text{ mod }2\pi\mathbb{Z}~,$ (2.6)
with a $\mathbb{Z}_{k}$ valued cocycle $\left(\omega_{D+1}\right)_{k}$. Then, its slant product $i_{g}[\left(\omega_{D+1}\right)_{k}]$ also has the value $2\pi\mathbb{Z}/k$. Now, suppose that the cohomology class $i_{g}\left(\omega_{D+1}\right)_{k}]$ has the order of $k^{\prime}$ in $H^{D}(BG,U(1))$. Then there exists a cocycle representative $(i_{g}\omega_{D+1})^{\prime}_{k^{\prime}}$ which takes value in $2\pi\mathbb{Z}/k^{\prime}$ for any $(D+1)$ group elements, and the cocycle $i_{g}[\left(\omega_{D+1}\right)_{k}]$ is written as
$i_{g}[\left(\omega_{D+1}\right)_{k}]=\frac{k}{k^{\prime}}\left(i_{g}\omega_{D+1}\right)^{\prime}_{k^{\prime}}+d(i^{\prime}_{g}\omega_{D+1})_{k^{\prime\prime}}/k^{\prime\prime}~,$ (2.7)
where the second term $d(i^{\prime}_{g}\omega_{D+1})_{k^{\prime\prime}}/k^{\prime\prime}$ represents the difference from a cocycle representative $\left(i_{g}\omega_{D+1}\right)^{\prime}_{k^{\prime}}$ by a coboundary. $(i^{\prime}_{g}\omega_{D+1})_{k^{\prime\prime}}$ takes value in $2\pi\mathbb{Z}/k^{\prime\prime}$ with some integer $k^{\prime\prime}$. We simply denote the above decomposition of $i_{g}\omega_{D+1}$ by
$i_{g}\omega_{D+1}=i^{A}_{g}\omega_{D+1}+\frac{1}{|\omega_{D+1}|}d(i^{B}_{g}\omega_{D+1})~,$ (2.8)
where $|\omega_{D+1}|=k$ is the order of $\omega_{D+1}]$ in $H^{D+1}(BG,U(1))$, and $i^{B}_{g}\omega_{D+1}\in\frac{2\pi}{k^{\prime\prime}}\mathbb{Z}$.
@@
The contribution from $i^{A}_{g}$ modifies the fusion rules of magnetic defects, while $i_{g}^{B}$ modifies the junction of magnetic defects [38] (see Figure 3). As a consequence, the fusion rule of the coset symmetry on the gapped boundary of the bulk $G$ gauge theory is also modified. The boundary condition breaks $G$ to subgroup $K$, and the boundary action for $K$ is $\alpha_{D}$. Below, we describe the fusion rule of coset symmetries realized by the topological defects of the boundary.
@@
$V_{g}\times V_{g^{\prime}}=V_{g+g^{\prime}}\frac{1}{{\cal N}}\left(\sum_{\lambda\in Z(G)}{\cal W}_{i_{\lambda}i_{g}\omega_{D+1}}\right)\left(\sum_{\lambda^{\prime}\in Z(G)}{\cal W}_{i_{\lambda^{\prime}}i_{g^{\prime}}\omega_{D+1}}\right)\big/\{{\cal W}_{i_{\lambda^{\prime\prime}}i_{g+g^{\prime}}(\omega_{D+1})}:\lambda^{\prime\prime}\in Z(G)\}~,$ (2.9)
@@
$V_{g}\times{\cal W}_{i_{\lambda}i_{g}\omega_{D+1}}=V_{g}~,\quad\forall\lambda\in Z(G)~,$ (2.10)
@@
Reflecting such fusion rules of the magnetic defects in the bulk, the fusion rules of coset symmetry on the boundary are modified if $\omega_{D+1}$ satisfies $\omega_{D+1}|_{K}=d\alpha$ and also $i_{g}^{A}\omega_{D+1}]_{K}\neq 0$. In other words, although the restriction of $\omega_{D+1}$ to subgroup $K$ is exact, it is no longer exact if we first take the slant product of $\omega_{D+1}$ with respect to $G$.
@@
The property that the coset symmetry is attached to a gauged SPT defect implies that the coset symmetry becomes “more non-invertible” in the presence of twists: for magnetic defects with center holonomy, there are non-invertible fusion rule obtained by replacing $i_{\lambda}i_{g}\omega_{D+1}$ in (2.9), (2.10) to $i_{\lambda}i_{g}\omega_{D+1}]_{K}$.
@@
Let us briefly recall how the slant product $i^{B}$ modifies the fusion rule of magnetic defects in the bulk $G$ gauge theory. The slant product $i^{B}$ modifies the (D2)(D-2)-dimensional junction of the magnetic defects by attaching the end of the gauged SPT defects in (D1)(D-1) dimensions. In other words, the fusion algebra of the magnetic defects is extended by the gauged SPT defects.
@@
$Z_{\omega}(G)\equiv\{g\in Z(G):\quad[i^{A}_{g}\omega_{D+1}]=0\in H^{D}(BG,U(1))\}~.$ (2.11)
@@
$1\rightarrow H^{D-1}(BG,U(1))\rightarrow{\cal G}^{(1)}\rightarrow Z_{\omega}(G)\rightarrow 1~,$ (2.12)
which is characterized by the second cohomology $\Omega_{\omega}$ [38]
$\Omega_{\omega}(g,g^{\prime})=\frac{1}{|\omega_{D+1}|}\left(i_{g}^{B}\omega_{D+1}+i_{g^{\prime}}^{B}\omega_{D+1}-i_{g+g^{\prime}}^{B}\omega_{D+1}\right)~,$ (2.13)
where $|\omega_{D+1}|$ is the order of $\omega_{D+1}]$ in the group $H^{D+1}(BG,U(1))$.
Let us now discuss how the above central extension affects the structure of the coset symmetry. We define $Z_{\omega}(G/K):=Z_{\omega}(G)/(Z_{\omega}(G)\cap K)$, which is the group of nontrivial coset symmetry defects on the boundary carrying the center holonomy. Here let us consider the cases where $Z_{\omega}(G)$ has the form of direct product, $Z_{\omega}(G)=Z_{\omega}(G/K)\times K^{\prime}$, with $K^{\prime}=Z_{\omega}(G)\cap K$. The generic cases will be studied in Sec. 2.3. In this case, $\Omega_{\omega}$ directly induces the central extension of the coset symmetry, and defines an extension class in $H^{2}(Z_{\omega}(G/K),H^{D-1}(BK,U(1)))$. This is obtained by restriction to the group $Z_{\omega}(G/K)\subset Z_{\omega}(G)$, together with the restriction of gauge group $G\to K$ in $H^{D-1}(BG,U(1))$.
The coset 0-form symmetry at the boundary then has the following central extension due to the bulk cocycle $\omega_{D+1}$,
$1\rightarrow H^{D-1}(BK,U(1))\rightarrow{\cal G}^{(0)}_{G/K}\rightarrow Z_{\omega}(G/K)\rightarrow 1~,$ (2.14)
characterized by the restriction of $\Omega_{\omega}$. As we will see in Sec. 2.3, in the presence of the boundary twist $\alpha_{D}$ the class for the central extension $\Omega_{\omega}$ will be further modified by slant product of $\alpha_{D}$.